f(x)=-2(x-1)^3(x+2)^2


1
Expert's answer
2021-05-07T11:49:02-0400

"f(X)=-2(x-1)^3(x+)^2"

"=-2(x^3-1^3-(3\times x\times 1)(x+1)(x^2+4+4x)"

"=-2(x^3-1-3x(x-1)(x^2+4+4x)"

"=[-2x^3+2+6x(x-1)](x2+4+4x)"

"=[-2x^3+2+6x^2-6x](x^2+4+4x)"

"=-2x^5-8x^3-8x^4+2x^2+8+8x+6x^4+24x^2+24x^3-6x^3-24x-24x^2"

"=-2x^5-2x^4+10x^3+2x^2-16x+8"